OpenStudy (anonymous):

How do I find the integral of sin(x)^2

OpenStudy (anonymous):

try integration by parts

OpenStudy (anonymous):

u = sinx dv = sinx du = cosx v = cosx sinxcosx - \[\int\limits_{?}^{?} \cos^2x\]

OpenStudy (anonymous):

i don't see how thats going to get me anywhere

OpenStudy (anonymous):

sorry is it \[\sin ^{2}x\] or \[\sin x ^{2}\]???

OpenStudy (anonymous):

the first.

OpenStudy (anonymous):

\[\sin ^2x\]

OpenStudy (anonymous):

\[\cos ^{2}x=1-\sin ^{2}x\] \[\int\limits_{?}^{?}\sin ^{2}xdx=-\sin x \cos x +\int\limits_{?}^{?}(1-\sin ^{2}x)dx\]

OpenStudy (anonymous):

therefore, the last integral of sin^2(x) can be put to the left hand side.

OpenStudy (anonymous):

giving integral (sin^2(x))=(-sin(x)cos(x)+x)/2

OpenStudy (anonymous):

would the two integrals divided by each other reduce to -1 though? so it'd be =sinxcosx +1 in the end?

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\sin^2x\] \[\div\] \[\int\limits_{}^{} 1-\sin^2x\] = -1?

OpenStudy (anonymous):

or is that just improper math? I'm not sure, I dont remember ever dividing integrals.

OpenStudy (anonymous):

oh wait, I wouldnt be dividng them though, I'd subtract, though I still think I'm missing something where does the 2 in -sinxcosx/2 come from?

OpenStudy (anonymous):

no, because you need to add integral (sin^2(x)) to both sides, not dividing them, giving 2 times integral (sin^2(x)) on the LHS and cancelling it on the RHS

OpenStudy (anonymous):

Oh, okay, I get it all now I forgot about separating integrals and what not. This all makes sense.

OpenStudy (anonymous):

My AP calc test is on wednesday, so, reviewing all this stuff is really helpful, but a pain in the retriceat the same time.

OpenStudy (anonymous):

good luck for wednesday.

OpenStudy (anonymous):

thanks.